OpenStudy (anonymous):

Check my work? Question: The variables x and y vary inversely. If x = 5/2 when y = 50, find an equation relating x and y. Answer: Okay so An equation of the form y = k/x expresses inverse variation and y has to equal 50 and x has to equal 5/2. So 50 = k/5/2. Now we multiply both sides by 2/5. So 50*2/5 = k/5/2*2/5 = 20 = k. y = 20x

OpenStudy (anonymous):

y=k/x 50 = k/(5/2) 50*5/2 =k k=125 So, y=125/x

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

Yes Thank you.

OpenStudy (anonymous):

welcome